17. Given the following energies in kcal/moles (CH3/CH3 eclipsed-4.1, CH3H eclipsed = 1.2, H/H eclipsed = 1, CHs/CH3 gauche 0.8), to the nearest 0.1 of a kcal/mole what is the energy difference between the following Wedge-Slash representations a. 0.8 b. 5.5 c. 6.3 d. 7.1 e. not a.-d
Answer #1

In the first conformation,

there are 2 gauche CH3/CH3 interactions ,

therefore total energy is equal to (0.8+0.8) Kcal/moles = 1.6Kcal/moles.

For the second conformation,

there are 1 CH3/CH3 eclipsed interactions, 1 CH3/H eclipsed, 1 H/H eclipsed and 1 CH3/CH3 gauche interactions ,

therefore total energy is equal to (4.1+1.2+1.0+0.8) = 7.1Kcal/moles

Energy difference = 7.1-1.6

= 5.5Kcal/moles